OpenStudy (anonymous):

given that the first 4 terms of a sequence are: 4, 12, 24, 40,.... find the nth term of this sequence.

OpenStudy (anonymous):

A1 = 4 n => 2 An = A(n-1) + 4n A2 = A1 + 4*2 = 4 + 8 = 12 A3 = A2 + 4*3 = 12 + 12 = 24 A4 = A3 + 4*4 = 24 + 16 = 40 ?

OpenStudy (anonymous):

omg. i, love youuuuu :D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i'm serious. without you, i would've died. :D:D:D:D